If three resistor are connected in series, what is the total current if the value of R1= 12 mega ohm R2 = 20 mega ohmR3 = 5 mega ohm and the voltage source is Et = 24 volts


1
Expert's answer
2021-04-01T01:10:57-0400

"R_1=12M \Omega=12 \times 10^6 \Omega"

"R_2=20M \Omega=20 \times 10^6 \Omega"

"R_3=5M \Omega=5 \times 10^6 \Omega"

"R_{Total}=(12+20+5) M \Omega =37M \Omega =37\times 10^6 \Omega"

The total current, "I= \frac{E_t}{R_{Total}}=\frac{24}{37\times 10^6}= 0.6486 \times10^{-6} A= 0.6486 mA"